OpenStudy (anonymous):

Someone please check. q+12-2(q-22)>0 q+10q-220>0 11q(-220+220)>0 11q>220 q>20

OpenStudy (anonymous):

where do you get 10q from?

OpenStudy (anonymous):

12-2, then I distributed by parenthesis

OpenStudy (anonymous):

Is that wrong NickR?

OpenStudy (anonymous):

you need to distribute first

OpenStudy (theeric):

\[q+12-2(q-22)>0\]

OpenStudy (theeric):

That's the start, right?

OpenStudy (theeric):

@cbm (just making a notification pop up :P )

OpenStudy (anonymous):

i think you made a mistake there q+12-2(q-22)>0 q+12-2q+44 >0 -q +56 >0 -q > -56 q>56

OpenStudy (anonymous):

sorry q<56

OpenStudy (theeric):

I agree with jayz657's work. Any questions, @cbm ?

OpenStudy (anonymous):

Thank ya'll so much! I was reworking it. Yep, once I distributed I got the same answer. The sign at the end changed right? From > to <

OpenStudy (anonymous):

I wish you could give more than one best response

OpenStudy (theeric):

Yep! Because you multiplied or divided by a negative!

OpenStudy (anonymous):

@theEric thanks!

OpenStudy (theeric):

My pleasure. I'm glad you got it worked out :) Good luck with any more you have to do!

OpenStudy (anonymous):

Thanks, I'm struggling with one more now trying to set up an inequality.

OpenStudy (theeric):

You can ask it here or in another question. I have time if you'd like a hand. Algebra like this can be hard at first.

OpenStudy (anonymous):

Ok, it will take me a minute to type it. Thanks for your help. Yes, I'm too old to learn this stuff.

OpenStudy (theeric):

Haha, no such thing. It might just require some thinking and rethinking, right? Maybe thinking in a different light! I started thinking about counting cookies/blocks again before so the numbers meant something.

OpenStudy (theeric):

And take your time!

OpenStudy (anonymous):

Your making a rectangular patio that is 10ft wide and 20ft long, using square pavers that have an area of 0.81 sq ft, what is the # of pavers u will need? Write and solve an inequality to find the answer.

OpenStudy (anonymous):

Lol, I'm trying. I thought I had equations but not the fractions

OpenStudy (anonymous):

would it be like (10x20).81=p

OpenStudy (anonymous):

idn

OpenStudy (theeric):

The first step is always understanding the problem. Then we move on.. So I have a quick question... Each "paver" is .81 square feet, and we need to know how many of the will fit in our 10 foot by 20 foot area?

OpenStudy (anonymous):

Yep, thats the question.

OpenStudy (theeric):

Alright! Then (10*20) is definately your area of the patio. And you wanna see how many .81 area portions will fit into that. How many times does .81 go into (10*20) is the question to get you to your answer. That's a division thing.

OpenStudy (anonymous):

I got 246,9 so P≥246.9

OpenStudy (theeric):

That's just one way to look at it. There's another way in which you build the equation and solve.. I got the same number after rounding. Is this supposed to be inequality?

OpenStudy (anonymous):

I'm not sure if I'm writing the inequality right or not (10ftx20ft)/.81=p

OpenStudy (anonymous):

Yes, it's supposed to be an inequality

OpenStudy (theeric):

That's what I would do to!

OpenStudy (anonymous):

To me it works and gives the answer when you work it.

OpenStudy (theeric):

I don't see why it would be inequality! There must be something I'm not thinking of.. I thought it would want an exact number...You want so many pavers.... OR! Do you want at least so many pavers? Meaning that many or more?

OpenStudy (anonymous):

Do you want me to create a problem so I can give you a medal? I'm not sure what the medal do.

OpenStudy (anonymous):

I was trying to figure out why it want come out to a whole number also unless it's ≥

OpenStudy (theeric):

They're just sort of recognition you can count. I'm not worried about it. Thanks though!

OpenStudy (theeric):

Yeah, I think maybe that's the idea of it.. why they want inequality.. Although you'd think rounding it up to 247 would be good enough! Maybe the question was just not appropriate for this subject!

OpenStudy (anonymous):

Oh, ok thanks! Maybe (10x20)/.81≥p

OpenStudy (theeric):

I would say \[p\ge 246.0\]

OpenStudy (theeric):

You would want more pavers rather than less. At least you get your patio done that way :P

OpenStudy (anonymous):

LOL, that's the truth.

OpenStudy (theeric):

Pavers area things should be greater than or equal to the patio area... to get the patio done :P Yeah!

OpenStudy (theeric):

Using just that sentence, you can set up the inequality. \[p_{NumberOfPavers}\ge A_{PatioArea}\]

OpenStudy (anonymous):

I've been trying everything to get an even number. I don't understand why it would be a decimal. I tried breaking it up into inches, etc. Thanks again for your help.

OpenStudy (theeric):

It might just not come out to be an even number. It happens. We could solve it to be a fraction instead of a decimal. With fractions, your always exact (no rounding).

OpenStudy (theeric):

I have five cookies to share between us. How many can we each have? 2.5. Books don't always make problems come out to nice numbers either..

OpenStudy (theeric):

So maybe that's the case here.

OpenStudy (anonymous):

true, I hope so.

OpenStudy (theeric):

Same. Good luck with that! If you solve it as a fraction, note that .81 = 81/100.

OpenStudy (anonymous):

alright, Thanks again!!

OpenStudy (theeric):

\[\frac{10*20}{.81}=\frac{10*20}{\frac{81}{100}}\] Np!

OpenStudy (anonymous):

:)

OpenStudy (theeric):

Just in case... From there, it's \[\frac{10*20*100}{81}=\frac{20000}{81}=246\frac{74}{81}\]

OpenStudy (anonymous):

Yep, that works good or at least it looks better. I figured I could write the equation as (p)x0.81≥200. I'm still working it to see if it will work.

OpenStudy (theeric):

That's great, and it was the other way I would think of the equation to solve!

OpenStudy (theeric):

And you don't have to change the sign because 0.81 is not negative.